Carol is 5 times as old as her brother she will be three times as old as her brother in 2 years.how old is Carol now?
Found 2 solutions by FrankM, macston:
Answer by FrankM(1040)

C=5B
C+2=3(B+2)
5B+2=3B+6
2B=4
B=2 C=10 Now
In 2 years, B=4 C=12



Answer by macston(5194)

.
B=brother's age; C=Carol's age=5B
.
C+2=3(B+2)
(5B)+2=3(B+2)
5B+2=3B+6
2B=4
B=2
.
Brother is 2 years old.
.
C=5B=5(2)=10
.
ANSWER: Carol is 10 years old.
.
CHECK:
.
C+2=3(B+2)
10+2=3(2+2)
12=3(4)
12=12
.
.